OpenStudy (anonymous):

help!?!?!? Find the volume of a cone with a diameter of 9 inches and a height that is 5 times the radius. Use 3.14 for pi and round your answer to the nearest hundredth. 47.1 in3 78.51in3 476.89 in3 1430.66 in3

OpenStudy (anonymous):



OpenStudy (anonymous):

replace all into the formula, you can get the answer

OpenStudy (anonymous):

1 over 3 (πr2)(h) 1 over 3 * 4.5 ^2*22.5 = 151.875

OpenStudy (anonymous):

@OOOPS 1 over 3 (πr2)(h) 1 over 3 * 4.5 ^2*22.5 = 151.875

OpenStudy (anonymous):

you have to find out r first because you don't have it. You have diameter, so radius = 1/2 diameter .

OpenStudy (anonymous):